# DGC-GNN: Descriptor-free Geometric-Color Graph Neural Network for 2D-3D Matching

### v1

@@ 3.2 Visual Localization
The visual localization task is the estimation of the 6 degrees-of-freedom camera pose of an input query image w.r.t a known map of the scene. One of the most prominent ways of approaching this problem is via establishing 2D-3D correspondences and running robust pose estimation. Following zhou2022geometry , we ran the proposed DGC-GNN to obtain matches. For each query image, we match its keypoints against the 3D points from the top-10 retrieved views to build the 2D-3D correspondences. The camera pose is then estimated by PnP-RANSAC fischler1981random ; 1217599 . We consider two standard datasets 7scenes shotton2013scene and Cambridge landmarks kendall2015posenet to evaluate the proposed method. For 7scenes, we extract the keypoints with the SIFT detectors, and the top 10 pairs are retrieved using DenseVLAD torii201524 . For Cambridge landmarks, the keypoints are extracted by SuperPoint detone2018superpoint to ensure consistency with the SuperPoint-based structure-from-motion model. The top 10 pairs are provided by NetVLAD Arandjelovic16 .
@@
Results. In Table 3, we present the 3D model maintenance costs, privacy, storage requirements, and camera pose median errors (cm, ${}^{\circ}$) of standard descriptor-based localization techniques and descriptor-free methods, including the proposed DGC-GNN. The proposed method consistently outperforms GoMatch on all scenes by a significant margin. On the Cambridge-Landmarks dataset, the average median pose error of DGC-GNN is 54 cm / ${}^{\circ}$, while GoMatch leads to 173 cm / ${}^{\circ}$ average error. On 7Scenes, the average error of DGC-GNN is 15 cm / ${}^{\circ}$ and that of GoMatch is 22 cm / ${}^{\circ}$. DGC-GNN requires a similar amount of memory to other descriptor-free methods. Also, it inherits their privacy-preserving properties due to not requiring and storing visual descriptors.
@@
The difference between descriptor-based (DB) and descriptor-free (DF) algorithms is clear. While descriptor-based ones lead to the best accuracy overall, they require excessive memory and descriptor maintenance and are susceptible to privacy attacks. Although the model compression method, HybridSC camposeco2019hybrid , shows effectiveness in storage saving, it achieves similar performance compared to DGC-GNN on the Cambridge dataset while still requiring descriptor maintenance. End-to-end methods (E2E) overcome these problems and achieve accurate results. However, their main limitation is that such approaches must be trained independently on each scene. The proposed DGC-GNN only needs to be trained once, making it more efficient and convenient to use as an off-the-shelf tool.
@@ 4 Conclusion
In conclusion, this paper introduces DGC-GNN, a novel graph-based pipeline for descriptor-free 2D-3D matching that effectively leverages geometric and color cues in a global-to-local manner. Our global-to-local procedure encodes both Euclidean and angular relations at a coarse level, forming a geometric embedding to guide the local point matching. By employing a cluster-based transformer, we enable efficient information passing within local clusters, ultimately leading to significant improvements in the number of correct matches and the accuracy of pose estimation. Compared to the state-of-the-art descriptor-free matcher GoMatch zhou2022geometry , the proposed DGC-GNN demonstrates a substantial improvement, doubling the accuracy on real-world and large-scale datasets. Furthermore, it results in significantly increased localization accuracy. These advancements contribute to reducing the gap between descriptor-based and descriptor-free methods while addressing the limitations of descriptor-based ones, such as memory footprint, maintenance costs, and susceptibility to privacy attacks. The source code and models will be made publicly available.
@@
Acknowledgements. This work was supported by the Academy of Finland (grant No. 327911, No. 353138), and by ETH Postdoc fellowship . We acknowledge the computational resources provided by the Aalto Science-IT project and CSC-IT Center for Science, Finland.
5 Supplementary Material
@@ Appendix A Training and Evaluation Details
Dataset Generation. The training data generation process for MegaDepth li2018megadepth follows the methodology outlined in GoMatch zhou2022geometry . The undistorted SfM model reconstructions used in MegaDepth are provided by D2Net dusmanu2019d2 . For training, we sample up to 500 images from each scene. For each sampled image, we select the top-$k$ co-visible views that have at least 35% image overlap. This ensures that there are enough matches available for training. The overlapping score is computed by dividing the number of co-visible 3D points by the total number of points in the training image.
@@
In the case of ScanNet dai2017scannet , a similar procedure is conducted. We also sample up to 500 images from each scene for the training set generation. The co-visible images are obtained using the co-visible scores provided by LOFTR sun2021loftr . We extract all the co-visible views of the training image with co-visible scores larger than 0.65, and then randomly sample the top-$k$ views for training. Since ScanNet is an RGB-D dataset without SfM models, we obtain the 3D points for each image by projecting the detected 2D keypoints with valid depth to the 3D coordinate. By doing this for each image, we reconstruct a sparse 3D point cloud based on the detected 2D keypoints. Note that the correspondence between different co-visible frames is not required in this case.
@@
In total, for MegaDepth, we generate a training set consisting of 25,624 images from 99 scenes and a test set comprising 12,399 images covering 53 scenes. For ScanNet, we create a training set with 52,008 images from 105 scenes. The test set for ScanNet consists of 14,892 query images from 30 scenes. The data generation of 7Scenes shotton2013scene and Cambridge Landmarks kendall2015posenet follows the same procedure in  zhou2022geometry .
@@
Inference. We consider a query with at least 10 keypoints as valid input. The 3D points from the top-$k$ retrieved database images are then applied to match against the queries with our proposed pipeline. We use the Sinkhorn algorithm sinkhorn1967concerning ; cuturi2013sinkhorn to optimize the extended cost matrix ¯$\bar{\mathcal{M}}\in\mathbb{R}^{N+1,M+1}$ in an iterative manner with up to 20 iterations to obtain the initial matches. The final matches are obtained by filtering the matches with matching confidence $\theta<0.5$ in the outlier rejection module. For the visual localization task, the camera poses are estimated by the P3P solver with RANSAC fischler1981random implemented in OpenCV bradski2000opencv and then refined by Levenberg-Marquardt more2006levenberg algorithm on the inliers matches, minimising the reprojection error.
@@
Figure 5: 2D-3D Matching (shown by green lines) with the proposed DGC-GNN and GoMatch zhou2022geometry .
Appendix B Additional Results
@@
Qualitative Results. More visualizations of inlier matches provided by DGC-GNN and GoMatch on MegaDepth are shown in Fig. 5. DGC-GNN consistently achieves more matches on multiple scenes, highlighting the effectiveness of the proposed method.
@@
Figure 7: Architectural Ablation Study. We plot the training loss and the number of matches estimated during the training for different architectures. Notably, we observe that the inclusion of color encoding in the training procedure leads to a significant improvement.
@@
Additional Ablation Results. In addition to the ablation results presented in the main paper, we also provide ablation results for single-view matching with $k=1$ on MegaDepth li2018megadepth . Furthermore, we conduct two additional ablations to investigate the impact of different component selections. Firstly, we compare the effectiveness of the geometric global embedding (G. Emb.) used in the main paper with the global clustering label embedding (G. Label). Instead of encoding geometric cues, we encode the label of each global cluster and concatenate it to the local point feature. Then, we explore the selection of different clustering algorithms. We compare the performance of K-Means and Mean-Shift clustering algorithms in our pipeline.
@@
The results are presented in Table 4. We observe similar conclusions for each component as in the main paper. The results obtained using the global label embedding (G. Label) with cluster attention (C.Att). show even worse performance compare to geometric embedding (G. Emb.) only, indicating the superiority of our clustering-based geometric embedding over the label embedding and highlighting the importance of incorporating geometric cues in the embedding process for effective point matching. Regarding the impact of different clustering algorithms, we only observe a minor difference in K-Means and Mean-Shift results, suggesting that our approach is robust to the choice of the clustering algorithm.
In addition to the numerical results, we also visualize the inlier matches (see Fig. 6), the training loss and the number of predicted matches during training (see Fig. 7) with different architectures to provide deep insights into the behaviour and performance of different architectures.
Appendix C Generalizability
Similar to zhou2022geometry , we discuss the generalizability of our DGC-GNN model on the visual localization task across different types of training and evaluation scenes. Specifically, we investigate the performance of our model when trained on MegaDepth li2018megadepth and ScanNet dai2017scannet , and evaluated on the 7Scenes dataset shotton2013scene . We also explore the impact of using different keypoint detectors, namely SIFT lowe2004distinctive and SuperPoint detone2018superpoint , during the evaluation.
The results of these experiments are summarized in Table 5, which provides an overview of the performance of our DGC-GNN model under different training and evaluation conditions. The results show that the DGC-GNN model trained on the ScanNet dataset, using the SIFT keypoint detector, and evaluated on the indoor 7Scenes dataset with the same detector, achieves the best overall performance. This demonstrates the effectiveness and robustness of our model in this specific training and evaluation scenario. However, we also note minor differences in performance when the model is trained on the MegaDepth dataset or evaluated using SuperPoint keypoints. Despite these variations, our findings suggest that the DGC-GNN model is scene-agnostic and can generalize well to different types of keypoint detectors. This is an encouraging result as showcases the versatility and generalization capability of the DGC-GNN approach, allowing it to be applied in various settings without significant degradation in performance.
Scenes Trained on MegaDepth li2018megadepth (SIFT) Trained on ScanNet dai2017scannet (SIFT)
GoMatch zhou2022geometry (SIFT) GoMatch zhou2022geometry (SP) DGC-GNN (SIFT) DGC-GNN (SP) DGC-GNN (SIFT) DGC-GNN (SP)
Chess 14 / 11.65 14 / 11.56 13 / 11.41 13 / 11.46 13 / 11.43 14 / 11.51
Fire 13 / 13.86 12 / 13.71 15 / 11.81 17 / 12.30 15 / 11.77 16 / 12.03
Heads 19 / 15.17 15 / 13.43 15 / 13.13 14 / 12.78 14 / 12.95 14 / 13.02
Office 11 / 12.48 17 / 11.76 17 / 11.66 17 / 11.66 16 / 11.61 17 / 11.66
Pumpkin 16 / 13.32 28 / 15.65 18 / 12.03 12 / 12.75 18 / 11.93 10 / 12.38
Redkitchen 13 / 12.84 14 / 13.03 18 / 12.14 10 / 12.36 18 / 12.09 19 / 12.28
Stairs 89 / 21.12 58 / 13.12 83 / 21.53 55 / 13.05 71 / 19.50 58 / 14.32
Table 5: Model Generalizability on Visual Localization Task. We report the translation and rotation median error (cm / ${}^{\circ}$) on 7Scenes dataset shotton2013scene . We evaluate the model trained on both MegaDepth li2018megadepth and ScanNet dai2017scannet datasets. The best performance is bold.
@@ Appendix D Model Parameters and Timing
We discuss the model parameters and running time of DGC-GNN in this section. DGC-GNN incorporate global geometric embedding and local clustering attention, which has around 5.7 million trainable parameters and an estimated model size of 22.6 MB. The average inference time for each image pair over the Megadepth evaluation queries is 77.8ms, and the number is 70.9ms on ScanNet queries. The measurements are conducted on a 32GB NVIDIA Telsa V100 GPU.

### v2

@@ 3.2 Visual Localization
Visual localization estimates the 6 degrees-of-freedom camera pose of an input query image w.r.t a known map of the scene. One of the most prominent ways of approaching this problem is via establishing 2D-3D correspondences and running robust pose estimation. Following [56], we ran the proposed DGC-GNN to obtain matches. For each query image, we match its keypoints with the 3D points from the top-10 retrieved views to build the 2D-3D correspondences. The camera pose is then estimated by PnP-RANSAC [16, 17]. We use two standard datasets, 7Scenes [46] and Cambridge Landmarks [22]. For 7Scenes, we extract the keypoints with the SIFT detectors, and the top 10 pairs are retrieved using DenseVLAD [50]. For Cambridge Landmarks, the keypoints are extracted by SuperPoint [11] to ensure consistency with the SuperPoint-based structure-from-motion model. The top 10 pairs are provided by NetVLAD [1].
3.2.1 Results
@@
In Table 3, we present the 3D model maintenance costs, privacy, storage requirements, and camera pose median errors (cm, ${}^{\circ}$) of standard descriptor-based localization techniques and descriptor-free methods. DGC-GNN consistently outperforms GoMatch on all scenes by a significant margin. On Cambridge Landmarks, the average median error of DGC-GNN is 54 cm / ${}^{\circ}$, while GoMatch leads to 173 cm / ${}^{\circ}$ error. On 7Scenes, the average error of DGC-GNN is 15 cm / ${}^{\circ}$, and that of GoMatch is 22 cm / ${}^{\circ}$. DGC-GNN requires a similar amount of memory to other descriptor-free methods. Also, it inherits their privacy-preserving properties due to not requiring visual descriptors.
@@
The trade-off between descriptor-based (DB) and descriptor-free (DF) algorithms is visible from the table. While descriptor-based ones lead to the best accuracy overall, they require excessive memory and descriptor maintenance and are susceptible to privacy attacks. Although the model compression method, HybridSC [6], shows effectiveness in storage saving, it achieves similar performance compared to DGC-GNN on the Cambridge Landmarks dataset while still requiring descriptor maintenance. End-to-end methods (E2E) overcome these problems and achieve accurate results. However, their main limitation is that such approaches must be trained independently on each scene. The proposed DGC-GNN only needs to be trained once, making it more efficient and convenient to use as an off-the-shelf tool.
3.2.2 Generalizability
Scenes Trained on MegaDepth [24] (SIFT) Trained on ScanNet [10] (SIFT)
GoMatch (SIFT) GoMatch (SP) DGC-GNN (SIFT) DGC-GNN (SP) DGC-GNN (SIFT) DGC-GNN (SP)
Chess 14 / 11.65 14 / 11.56 13 / 11.41 13 / 11.46 13 / 11.43 14 / 11.51
Fire 13 / 13.86 12 / 13.71 15 / 11.81 17 / 12.30 15 / 11.77 16 / 12.03
Heads 19 / 15.17 15 / 13.43 15 / 13.13 14 / 12.78 14 / 12.95 14 / 13.02
Office 11 / 12.48 17 / 11.76 17 / 11.66 17 / 11.66 16 / 11.61 17 / 11.66
Pumpkin 16 / 13.32 28 / 15.65 18 / 12.03 12 / 12.75 18 / 11.93 10 / 12.38
Redkitchen 13 / 12.84 14 / 13.03 18 / 12.14 10 / 12.36 18 / 12.09 19 / 12.28
Stairs 89 / 21.12 58 / 13.12 83 / 21.53 55 / 13.05 71 / 19.50 58 / 14.32
All 22 / 15.78 18 / 14.61 17 / 14.82 14 / 13.77 15 / 14.47 14 / 13.89
Table 4: Model Generalizability on Visual Localization Task. We report the translation and rotation median error (cm / ${}^{\circ}$) on 7Scenes dataset [46]. We evaluate the models trained on MegaDepth [24] and ScanNet [10] datasets. The best performance is bold.
Similar to [56], we discuss the generalizability of our DGC-GNN model on the visual localization task across different training and evaluation scenes. Specifically, we investigate the performance of our model when trained on MegaDepth [24] and ScanNet [10] and evaluated on the 7Scenes dataset [46]. We also explore the impact of using different keypoint detectors, namely SIFT [27] and SuperPoint [11], during the evaluation.
These experiments are summarized in Table 4, providing an overview of the performance of DGC-GNN under different training and evaluation conditions. While the best overall performance is achieved by training on the MegaDepth dataset with SIFT features, the results are similar in both training scenarios, showcasing that the proposed method generalizes well to unseen data. While we train on SIFT features, the best results are achieved by using SuperPoint features at inference time. This demonstrates that DGC-GNN is insensitive to the features used and can be utilized off-the-shelf even without retraining to our specific scenario.
@@ 4 Conclusion
In conclusion, this paper introduces DGC-GNN, a novel graph-based pipeline for visual descriptor-free 2D-3D matching that effectively leverages geometric and color cues in a global-to-local manner. Our global-to-local procedure encodes both Euclidean and angular relations at a coarse level, forming a geometric embedding to guide the local point matching. By employing a cluster-based transformer, we enable efficient information passing within local clusters, ultimately leading to significant improvements in the number of correct matches and the accuracy of pose estimation. Compared to the state-of-the-art descriptor-free matcher GoMatch [56], the proposed DGC-GNN demonstrates a substantial improvement, doubling the accuracy on real-world and large-scale datasets. Furthermore, it results in significantly increased localization accuracy. These advancements contribute to reducing the gap between descriptor-based and descriptor-free methods while addressing the limitations of descriptor-based ones, such as memory footprint, maintenance costs, and susceptibility to privacy attacks.
@@
Acknowledgements. This work was supported by the Academy of Finland (grants No. 327911, No. 353138) and the Hasler Stiftung Research Grant via the ETH Zurich Foundation. We acknowledge the computational resources provided by the CSC-IT Center for Science, Finland.
Supplementary Material
@@ Appendix A Training and Evaluation Details
Dataset Generation. The training data generation process for MegaDepth [24] follows the methodology outlined in GoMatch [56]. The undistorted SfM model reconstructions used in MegaDepth are provided by D2Net [14]. For training, we sample up to 500 images from each scene. For each sampled image, we select the top-$k$ co-visible views that have at least 35% image overlap. This ensures that there are enough matches for training. The overlapping score is computed by dividing the number of co-visible 3D points by the total number of points in the training image.
@@
In the case of ScanNet [10], a similar procedure is conducted. We also sample up to 500 images from each scene for the training set generation. The co-visible images are obtained using the co-visible scores provided by LoFTR [49]. We extract all the co-visible views of a training image with co-visible scores larger than 0.65. Then, we randomly sample the top-$k$ views for training. Since ScanNet is an RGB-D dataset without an SfM reconstruction, we obtain the 3D points for each image by projecting the detected 2D keypoints with valid depth to 3D. By doing this for each image, we reconstruct a sparse 3D point cloud based on the detected 2D keypoints. Note that the correspondence between different co-visible frames is not required in this case.
@@
In total, for MegaDepth, we generate a training set consisting of 25,624 images from 99 scenes and a test set comprising 12,399 images covering 53 scenes. For ScanNet, we create a training set with 52,008 images from 105 scenes. The test set for ScanNet consists of 14,892 query images from 30 scenes. The data generation of 7Scenes [46] and Cambridge dataset [22] follows the same procedure in  [56].
Figure 5: Points Reprojection and Image Recovery Example.
Method Reprojection InvSFM [35]
Points Points+RGB Points Points+RGB Points+SIFT
SSIM ($\downarrow$) 0.240 0.258 0.352 0.375 0.476
Table 5: SSIM Results. We evaluate the SSIM from Point Reprojection and Image Recovering, adding RGB to points leads only to a slight SSIM increase on both reprojection and image recovery.
@@
Inference. We consider a query with at least 10 keypoints as valid input. The 3D points from the top-$k$ retrieved database images are then applied to match against the queries with our proposed pipeline. We use the Sinkhorn algorithm [47, 9] to optimize the extended cost matrix ¯$\bar{\mathcal{M}}\in\mathbb{R}^{N+1,M+1}$ in an iterative manner with up to 20 iterations to obtain the initial matches. The final matches are obtained by filtering the matches with matching confidence $\theta<0.5$ in the outlier rejection module. For the visual localization task, the camera poses are estimated by the P3P solver with RANSAC [16] implemented in OpenCV [4] and then refined by Levenberg-Marquardt [29] algorithm on the inliers matches, minimizing the reprojection error.
@@
Figure 6: 2D-3D Matching (shown by green lines) with the proposed DGC-GNN and GoMatch [56].
Appendix B Privacy Issue of RGB Points
We investigate the impact on privacy resulting from incorporating RGB information into pixels and points. To assess this, we compute the Structural Similarity Index Measure (SSIM) for 3D points reprojected onto the image plane against the ground truth (GT) images on MegaDepth over 500 images from multiple scenes. Additionally, we recover the images from points + RGB and points + descriptors with InvSFM [35] to calculate the SSIM against the GTs. The findings are detailed in Table 5 and Fig 5. The addition of RGB data to points results in only a marginal increase in SSIM for both direct reprojection and image reconstruction via InvSFM, significantly less than what is achieved by incorporating SIFT descriptors. It is worth noting that denser point clouds might provide sufficient context, potentially leading to privacy concerns. However, in our setting, we mitigate this risk by limiting the number of keypoints from each database image to a maximum of 1024.
Appendix C Additional Results
@@
Qualitative Results. More visualizations of inlier matches provided by DGC-GNN and GoMatch on MegaDepth are shown in Fig. 6. DGC-GNN consistently finds more correct matches on multiple scenes, highlighting the effectiveness of the proposed method.
@@
Additional Ablation Results. In addition to the ablation results presented in the main paper, we also provide ablation results for single-view matching with $k=1$ on MegaDepth [24]. Furthermore, we conduct two additional ablations to investigate the impact of different component selections. Firstly, we compare the effectiveness of the geometric global embedding (G. Emb.) used in the main paper with the global clustering label embedding (G. Label). Instead of encoding geometric cues, we encode the label of each global cluster and concatenate it to the local point feature. Then, we explore the selection of different clustering algorithms. We compare the performance of K-Means and Mean-Shift clustering algorithms in our pipeline. Last, we study the effectiveness of the outlier rejection (OR) network.
@@
The results are presented in Table 6. We observe similar conclusions for each component as in the main paper. The results obtained using the global label embedding (G. Label) with cluster attention (C.Att) show even worse performance compared to geometric embedding (G. Emb.) only, indicating the superiority of our clustering-based geometric embedding over the label embedding and highlighting the importance of incorporating geometric cues in the embedding process for effective point matching. Regarding the impact of different clustering algorithms, we only observe a minor difference in K-Means and Mean-Shift results, suggesting that our approach is robust to the choice of the clustering algorithm. The results also demonstrate that outlier rejection is an essential post-processing module to achieve good performance. In addition to the numerical results, we visualize the inlier matches (see Fig. 7) to provide deep insights into the behavior and performance of different architectures.
Hyperparameters analysis. Besides the component ablations, we also give an in-depth analysis of the hyperparameters used in our main pipeline. Here, we add additional ablations on the number of input keypoints, the number of cluster groups at the coarse level, the number of nearest neighbors in the local graph build, and the outlier rejection threshold by retraining our DGC-GNN. The results are presented in Table 7. We observe that DGC-GNN with G. Clusters = 10 and Local NN = 10 achieves overall the best performance. Setting the outlier rejection threshold to 0.7 leads to the best performance. However, the results are stable across different configurations, indicating robustness to the parameter setting.
Methods G. Cluters Local NN OR Threshold Reproj. AUC (%) Rotation (${}^{\circ}$) Translation
@1 / 5 / 10px ($\uparrow$) Quantile@25 / 50 / 75% ($\downarrow$)
DGC-GNN 10 10 0.5 15.30 / 51.70 / 60.01 0.07 / 0.26 / 15.41 0.01 / 0.02 / 0.57
HyperParam. 5 10 0.5 14.73 / 50.12 / 58.26 0.08 / 0.28 / 18.76 0.01 / 0.03 / 0.99
15 10 0.5 15.14 / 50.56 / 58.62 0.07 / 0.28 / 17.66 0.01 / 0.03 / 0.89
10 20 0.5 14.77 / 49.84 / 57.97 0.07 / 0.29 / 18.26 0.01 / 0.03 / 0.90
10 30 0.5 14.75 / 50.95 / 59.45 0.08 / 0.28 / 15.48 0.01 / 0.03 / 0.58
10 10 0.3 13.28 / 46.44 / 55.05 0.08 / 0.43 / 18.63 0.01 / 0.04 / 0.98
10 10 0.7 16.63 / 56.26 / 64.46 0.07 / 0.19 / 12.58 0.01 / 0.02 / 0.27
Table 7: Ablation Study on Hyperparameters. We report the results of ablations with retrieved image $k=10$ on the number of global clusters, the number of nearest neighbour points for local graph build, and different thresholds for outlier rejection. The best results are bold.
Matching Results in pixel threshold. As mentioned in the main paper, we selected the ground truth matches in normalized image coordinates. The described GT difference only affects the reprojection AUC scores. Here, we present the matching results in Table 5 by selecting the ground truth matches in pixel coordinate with $1$ pixel threshold as done in [56]. Our conclusions still hold.
@@ Appendix D Model Parameters and Timing
We discuss the model parameters and running time of DGC-GNN in this section. DGC-GNN incorporates global geometric embedding and local clustering attention, which has around 5.7 million trainable parameters and an estimated model size of 22.6 MB. The average inference time for each image pair over the Megadepth evaluation queries is 77.8ms. It roughly breaks down into point encoding (24 ms), global geometric embedding (14 ms), cluster-based attention (22 ms), optimal transport (7 ms), and outlier rejection (8 ms). The measurements are conducted on a 32GB NVIDIA Telsa V100 GPU with a maximum of 1024 keypoints.